OpenStudy (studyguy2019tim):

Consider the functions f(x) = 3x − 2 and g(x) = x^2 − 1. What is the value of f[g(5)]? 312 70 168 −1 Confused! Please help with how to solve!

OpenStudy (marcelie):

plugin 5 into g (x)

OpenStudy (marcelie):

after that plugin ur answer what u got from g(x) then plug that into f(x)

OpenStudy (marcelie):

\[g (5) = (5)^2 -1 = 24\]

OpenStudy (marcelie):

\[f(24)= 3( 24)-2 = 72-2= 70\]

OpenStudy (studyguy2019tim):

That's what I thought! Your so smart! Thank you!

OpenStudy (marcelie):

no problem :)